# Random Utility with Aggregation

### v2

@@ 5.2.2 Maximum and minimum biases and the independent case
To construct Figure 1, we assumed $\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$. In this section, we relax this assumption and vary the values of $\lambda_{\{y,a_{0}\}}$ and $\lambda_{\{x,y,a_{0}\}}$ and report the resulting maximum and minimum biases in Figure 2. Each point in the figure corresponds to four possible bias measures, evaluated at nonnegative values of $\lambda_{\{x,y,a_{0}\}}(\{w\})$ and $\lambda_{\{x,y,a_{0}\}}(\{z\})$ satisfying $\lambda_{\{x,y,a_{0}\}}(\{w\})+\lambda_{\{x,y,a_{0}\}}(\{z\})\leq 1$:
  • (1) Maximum positive bias, obtained by varying all values of $\lambda_{\{x,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$ (red)
  • (2) Maximum negative bias, obtained by varying all values of $\lambda_{\{x,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$ (purple)
  • (3) Minimum absolute bias, obtained by varying all values of $\lambda_{\{x,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$ (blue)
@@
Figure 2: Maximum positive bias (red), maximum negative bias (purple), minimum absolute bias (blue), and bias under independent $\lambda$ (green), plotted across values of $\lambda_{\{x,y,a_{0}\}}$. The horizontal axis reports $\lambda_{\{x,y,a_{0}\}}(\{w\})$ and the vertical axis reports $\lambda_{\{x,y,a_{0}\}}(\{z\})$. The maximum and minimum biases are obtained by optimizing over all admissible values of $\lambda_{\{x,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$.
@@
The figure illustrates that, in the independent case, the biases are significantly smaller than the maximum biases and are often close to the minimum biases. [^17] In fact, when $\lambda_{\{x,y,a_{0}\}}(\{w\})=1$ or $\lambda_{\{x,y,a_{0}\}}(\{w\})=0$, the blue points and the green points coincide. Notably, both the maximum positive and negative biases can be substantial in magnitude. Specifically, the bias can exceed $2$ or fall below 3-3. As mentioned, when the bias $(\hat{u}(x)-\hat{u}(y))-(u(x)-u(y))$ is less than 1-1, it implies that the estimated utilities satisfy $\hat{u}(x)<\hat{u}(y)$, despite the true ordering being $u(x)>u(y)$. Moreover, the difference is large enough that $y$ appears substantially better than $x$ in the estimates. This demonstrates that bias distorts not only the estimated utility levels but also the inferred preference orderings, potentially leading to incorrect conclusions and decisions.
Conversely, when the bias exceeds $2$, we obtain $\hat{u}(x)-\hat{u}(y)>3$, indicating a substantial overstatement of the utility difference. In terms of odds ratios, this implies
$\frac{\exp(\hat{u}(x))}{\exp(\hat{u}(y))}\;>\;e^{2}\cdot\frac{\exp(u(x))}{\exp(u(y))}.$
@@
In other words, the estimated odds ratio overstates the true odds ratio by a factor greater than $e^{2}\approx 7$. Such distortions in relative odds can result in highly inaccurate predictions of choice probabilities.
5.2.3 Effect of preference structure on bias
In the last simulation on biases, we vary the preference structure to evaluate how overlapping preferences affect bias. We assume that the true unobservable dataset $\rho_{\scriptscriptstyle\mathscr{X}}$ is generated by logit models with $u(x)=2$, $u(y)=1$, and various values of $u(z)$ and $u(w)$ on a uniform grid over [5,5][-5,5]. The composition distributions $\lambda$ are fixed to be non-independent, since $\lambda$ is not the focus of this analysis. Specifically, we set $\lambda_{\{x,y,a_{0}\}}=(0.8,0.1,0.1)$, $\lambda_{\{x,a_{0}\}}=(0.8,0.1,0.1)$, and $\lambda_{\{y,a_{0}\}}=(0.1,0.8,0.1)$, where the three coordinates represent the probabilities of $\{z\}$, $\{w\}$, and $\{z,w\}$, respectively.
Figure 3 displays the resulting heatmap, with $u(z)$ and $u(w)$ on the axes. Red indicates positive bias at the top left corner and blue indicates negative bias at the bottom right corner, with color intensity reflecting the magnitude. The figure confirms our Conjecture (A): bias is largest when the preference structure is highly overlapping. In particular, as $u(z)$ and $u(w)$ move farther apart, cells lie farther from the diagonal, and the magnitude of bias increases. Conversely, when both $u(z)$ and $u(w)$ are either very high or very low, the bias decreases.
Figure 3: Heatmap of biases across values of $u(z)$ and $u(w)$, with $u(x)=2$, $u(y)=1$, and composition distributions fixed at $\lambda_{\{x,y,a_{0}\}}=(0.8,0.1,0.1)$, $\lambda_{\{x,a_{0}\}}=(0.8,0.1,0.1)$, and $\lambda_{\{y,a_{0}\}}=(0.1,0.8,0.1)$. The horizontal axis reports $u(z)$, the vertical axis reports $u(w)$, and color intensity indicates the magnitude of the bias defined in (9) (red for positive bias, blue for negative bias). Biases are largest when preferences overlap strongly, i.e., when $u(z)$ and $u(w)$ straddle $u(x)=2$ and $u(y)=1$.
@@ 5.3 Distance to the ARU Polytope
The previous subsection confirmed Conjecture (A), namely that estimation biases increase as the composition distribution $\lambda$ becomes more menu dependent and the preference structure becomes more overlapping. The purpose of this subsection is to explain why it holds. Our explanation is provided by Conjecture (B), which states that the distance from the observed dataset $\rho$ to the ARU polytope grows as the composition distribution $\lambda$ becomes more menu dependent or as the underlying preferences become more overlapping.
To quantify the deviation of the observable dataset $\rho$ from ARU-rationality, we calculate the minimum distance between $\rho$ and the set $\mathcal{P}_{\mathscr{A}}$ of ARUMs (i.e., the ARU polytope). (See Definition 3.7 in Section 3.1 for the definition.) Recall that a dataset $\rho$ belongs to $\mathcal{P}_{\mathscr{A}}$ if and only if it is ARU rational. Hence, the minimal distance from $\rho$ to $\mathcal{P}_{\mathscr{A}}$ provides a natural metric for assessing the extent of violation of the ARU rationality. The minimum distance is defined as follows:
$\min_{\rho^{\prime}\in{\mathcal{P}}_{\mathscr{A}}}\|\rho-\rho^{\prime}\|.$ (10)
@@
To see how the minimum distance can be calculated in our setup, remember that we have three alternatives $x,y,a_{0}$. Thus the number of linear orders is $6$; the number of coordinates of $\rho$ (i.e., $(D,i)$ such that $i\in D\subseteq\{x,y,a_{0}\}$) is 12. Then, the distance (10) can be calculated as follows: given $\rho\in\mathbf{R}^{12}$,
$\min_{\mu\in\Delta^{5}}\left\|\rho-\sum_{i=1}^{6}\mu_{i}\rho^{\succ_{i}}_{\mathscr{D}}\right\|_{2}^{2},$ (11)
where the norm is the Euclidean norm and $\rho^{\succ_{i}}_{\mathscr{D}}$ is defined by equation (6). Here, $\Delta^{5}$ is the five dimensional simplex and an element of the simplex satisfies $\mu\in\mathbb{R}^{6}$, $\mu_{i}\geq 0$, and $\sum_{i}\mu_{i}=1$.
@@ 5.3.1 Effect of composition distribution $\lambda$ on distance
Figure 5: Heatmap of the distance defined in (11) across values of $\lambda_{\{x,a_{0}\}}$, with $\lambda_{\{y,a_{0}\}}$ and $\lambda_{\{x,y,a_{0}\}}$ fixed at $(0.8,0.1,0.1)$ and utilities fixed at $u(x)=2$, $u(y)=1$, $u(z)=3$, and $u(w)=0$. The horizontal axis reports $\lambda_{\{x,a_{0}\}}(\{z\})$, the vertical axis reports $\lambda_{\{x,a_{0}\}}(\{w\})$, and color intensity indicates the magnitude of the distance. The blue-outlined cell marks the independent case ($\lambda_{\{x,a_{0}\}}=\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$); cells farther from the blue cell correspond to more menu-dependent cases, where the distance is larger.
@@
We compute the distance at each grid point. Figure 5 reports the resulting heatmap. Consistent with Proposition 4.5, the distance is zero in the independent case (blue-outlined cell). As $\lambda$ moves farther from the menu independence, the distance increases, confirming Conjecture (B).
To understand the figures, recall that we consider three alternatives $x,y$, and $a_{0}$. Thus, it can be shown that $\rho$ is ARUM if and only if it satisfies the standard monotonicity condition for all menus. Since $\rho$ is RU-rational, it satisfies Limited Monotonicity, that is monotonicity on $\{x,y\}$ with respect to adding $a_{0}$. Also, since $\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$, monotonicity is satisfied on menu $\{y,a_{0}\}$ with respect to adding $x$. Thus the only possible ways to violate monotonicity are (i) $\rho(\{x,y,a_{0}\},a_{0})>\rho(\{x,a_{0}\},a_{0})$ and (ii) $\rho(\{x,y,a_{0}\},a_{0})>\rho(\{y,a_{0}\},a_{0})$. The first occurs when $a_{0}$ is less attractive on $\{x,a_{0}\}$ than on $\{x,y,a_{0}\}$ and the second occurs when $a_{0}$ is more attractive on $\{x,a_{0}\}$ than on $\{x,y,a_{0}\}$. Since $u(z)$ is high and $u(w)$ is low, the first case occurs when $a_{0}$ is more likely to be $\{w\}$ (e.g., $\lambda_{\{x,a_{0}\}}(\{w\})=1$); the second case occurs when $a_{0}$ is more likely to be $\{z\}$ or $\{z,w\}$ (e.g., $\lambda_{\{x,a_{0}\}}(\{w\})=0$). These are two cases where the distance is large in Figure 5.
@@ 5.3.2 Effect of preference structure on distance
We produced a heatmap with axes corresponding to $u(z)$ and $u(w)$ and color intensity represents the magnitude of the distance. Figure 6 shows that the distance is largest in the lower-right corner, where the preference structure exhibits strong overlap, and smallest along the diagonal, where preferences are closer to non-overlapping, which is consistent with Proposition 4.2 and Conjecture (B).
Figure 6: Heatmap of the distance defined in (11) across values of $u(z)$ and $u(w)$, with $u(x)=2$, $u(y)=1$, and composition distributions fixed at $\lambda_{\{x,y,a_{0}\}}=(0.8,0.1,0.1)$, $\lambda_{\{x,a_{0}\}}=(0.8,0.1,0.1)$, and $\lambda_{\{y,a_{0}\}}=(0.1,0.8,0.1)$. The horizontal axis reports $u(z)$, the vertical axis reports $u(w)$, and color intensity indicates the magnitude of the distance. The distance is largest in the lower-right corner and smaller along the diagonal.
@@ 6 Summary
We study discrete choice with aggregated categories whose compositions are unobserved and show that RU-rationality imposes only limited implications for the observed dataset $\rho$: (i) Limited Monotonicty (adding aggregated categories to a menu of non-aggregated categories decreases choice probabilites) (ii) Partial RU-rationality (standard RUM consistency on menus that do not contain aggregated categories). We also characterize the admissible behaviors of consumers under RU-rationality and show that the characterization includes various irrational behaviors under ARU-rationality.
These results show how precisely RU-rationality is weaker than ARU-rationality, which is commonly assumed in empirical work. Given this, we identify two conditions under which RU-rationality implies ARU-rationality: (i) menu-independence of $\lambda$ and (ii) a non-overlapping preference structure. The following figure summarizes Theorem 3.1:
ARURemark 2.7
+ menu independent λ or

### v3

@@ 5.2.3 Maximum and minimum biases and the independent case
[…]
(1) Maximum positive bias, obtained by varying $\lambda_{\{x,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$ (red)
(2) Maximum negative bias, obtained by varying $\lambda_{\{x,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$ (purple)
(3) Minimum absolute bias, obtained by varying $\lambda_{\{x,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$ (blue)
@@
Figure 3: Maximum positive bias (red), maximum negative bias (purple), minimum absolute bias (blue), and bias under menu-independent $\lambda$ (green), plotted across values of $\lambda_{\{x,y,a_{0}\}}$. The horizontal axis reports $\lambda_{\{x,y,a_{0}\}}(\{w\})$ and the vertical axis reports $\lambda_{\{x,y,a_{0}\}}(\{z\})$. The maximum and minimum biases are obtained by optimizing over all admissible values of $\lambda_{\{x,a_{0}\}}$ and $\lambda_{\{y,a_{0}\}}$.
@@
The figure illustrates that, in the menu-independent case, the biases are significantly smaller than the maximum biases and are often close to the minimum biases. [^24] In fact, when $\lambda_{\{x,y,a_{0}\}}(\{w\})=1$ or $\lambda_{\{x,y,a_{0}\}}(\{w\})=0$, the blue points and the green points coincide. Notably, both the maximum positive and negative biases can be substantial in magnitude. Specifically, the bias can exceed $2$ or fall below 3-3. As mentioned, when the bias $(\hat{u}(x)-\hat{u}(y))-(u(x)-u(y))$ is less than 1-1, it implies that the estimated utilities satisfy $\hat{u}(x)<\hat{u}(y)$, despite the true ordering being $u(x)>u(y)$. Moreover, the difference is large enough that $y$ appears substantially better than $x$ in the estimates. This demonstrates that bias distorts not only the estimated utility levels but also the inferred preference orderings, potentially leading to incorrect conclusions and decisions.
@@
Conversely, when the bias exceeds $2$, we have $\frac{\exp(\hat{u}(x))}{\exp(\hat{u}(y))}\;>\;e^{2}\cdot\frac{\exp(u(x))}{\exp(u(y))}$. This means the estimated odds ratio overstates the true odds ratio by a factor greater than $e^{2}\approx 7$. Such distortions in relative odds can result in highly inaccurate predictions of choice probabilities.
5.2.4 Biases in parametric setup
In the previous sections, we treat each alternative such as $x$ as a label. Some readers may wonder how the results change if we consider covariates vector $x$, where utility is given by $\beta\cdot x$. In general, the results would be similar and large biases would be observed. This is because any bias on the utility level would be reflected in the estimate of $\beta$. For example, suppose each alternative is a vector containing price and quality. Assume that $x$ is higher quality and more expensive, and $y$ is lower quality and cheaper. As in the previous sections, assume that the true utility of $x$ is higher than that of $y$, this indicates that the coefficient $\beta_{quality}$ is larger than the coefficient $\beta_{price}$ in the true utility function.
The previous sections show that when $\lambda$ is significantly menu-dependent and $\mu_{\mathscr{X}}$ is significantly overlapping, then the estimated utility of $y$ may become larger than that of $x$. This implies that the estimated $\hat{\beta}_{quality}$ is smaller than $\hat{\beta}_{price}$. In this way, we expect that whenever the utility estimate is biased, we should observe similar biases in parameter estimation.
@@ 5.3 Distance to the ARU polytope
The previous subsections confirmed that estimation biases increase as the composition distribution $\lambda$ becomes more menu dependent and the preference structure becomes more overlapping. The purpose of this subsection is to explain why it holds. We observe that the distance from the reduced dataset $\rho$ to the ARU polytope grows as the composition distribution $\lambda$ becomes more menu dependent or as the underlying preferences become more overlapping.
To quantify the deviation of $\rho$ from ARU-rationality, we calculate the minimum distance between $\rho$ and the set $\mathcal{P}_{\mathscr{A}}$ of ARUMs (i.e., the ARU polytope):
$\min_{\rho^{\prime}\in{\mathcal{P}}_{\mathscr{A}}}\|\rho-\rho^{\prime}\|^{2}.$ (8)
See Definition 3.1 for the definition of the ARU polytope. Recall that a dataset $\rho$ belongs to $\mathcal{P}_{\mathscr{A}}$ if and only if it is ARU rational. Hence, the minimal distance from $\rho$ to $\mathcal{P}_{\mathscr{A}}$ provides a natural metric for assessing the extent of violation of the ARU rationality.
@@
To see how the minimum distance can be calculated in our setup, remember that we have three alternatives $x,y,a_{0}$. Thus the number of linear orders is six; the number of coordinates of $\rho$ (i.e., $(D,i)$ such that $i\in D\subseteq\{x,y,a_{0}\}$) is twelve. Then, the distance (8) can be calculated as follows: given $\rho\in\mathbf{R}^{12}$,
$\min_{\mu\in\Delta^{5}}\Big\|\rho-\sum_{i=1}^{6}\mu_{i}\rho^{\succ_{i}}_{\mathscr{D}}\Big\|^{2},$ (9)
where $\rho^{\succ_{i}}_{\mathscr{D}}$ is defined by equation (5). Here, $\Delta^{5}$ is the five dimensional simplex and an element of the simplex satisfies $\mu\in\mathbf{R}^{6}$, $\mu_{i}\geq 0$, and $\sum_{i}\mu_{i}=1$.
@@ 5.3.1 Effect of composition distribution $\lambda$ on distance
a Heatmap of the distance defined in (9) across values of $\lambda_{\{x,a_{0}\}}$, with $\lambda_{\{y,a_{0}\}}$ and $\lambda_{\{x,y,a_{0}\}}$ fixed at $(0.8,0.1,0.1)$ and utilities fixed at $u(x)=2$, $u(y)=1$, $u(z)=3$, and $u(w)=0$. The horizontal axis reports $\lambda_{\{x,a_{0}\}}(\{z\})$, the vertical axis reports $\lambda_{\{x,a_{0}\}}(\{w\})$, and color intensity indicates the magnitude of the distance. The blue-outlined cell marks the independent case ($\lambda_{\{x,a_{0}\}}=\lambda_{\{y,a_{0}\}}=\lambda_{\{x,y,a_{0}\}}$); cells farther from the blue cell correspond to more menu-dependent cases, where the distance is larger.
b Heatmap of the distance defined in (9) across values of $u(z)$ and $u(w)$, with $u(x)=2$, $u(y)=1$, and composition distributions fixed at $\lambda_{\{x,y,a_{0}\}}=(0.8,0.1,0.1)$, $\lambda_{\{x,a_{0}\}}=(0.8,0.1,0.1)$, and $\lambda_{\{y,a_{0}\}}=(0.1,0.8,0.1)$. The horizontal axis reports $u(z)$, the vertical axis reports $u(w)$, and color intensity indicates the magnitude of the distance. The distance is largest in the lower-right corner and smaller along the diagonal.
@@
We compute the distance at each grid point. Figure 5a reports the resulting heatmap. Consistent with Proposition 4.2, the distance is zero in the menu-independent case (blue-outlined cell). As $\lambda$ moves farther from the menu independence, the distance increases. We discuss the intuition underlying Figure 5a in Section E.1.
@@ 5.3.2 Effect of preference structure on distance
We produced a heatmap with axes corresponding to $u(z)$ and $u(w)$ and color intensity represents the magnitude of the distance. Figure 5b shows that the distance is largest in the lower-right corner, where the preference structure exhibits strong overlap, and smallest along the diagonal, where preferences are closer to non-overlapping, which is consistent with Proposition 4.1. Section E.1 in the online appendix provides further intuition for this figure.
@@ 6 Summary
We study stochastic choice over aggregates when the analyst does not observe the underlying alternatives each aggregate represents and the composition can vary across menus. RUM defined over underlying alternatives describes consumer’s actual choice among the alternatives available to them instead of analyst-defined aggregates. An ARUM, by contrast, is an empirical tool that enables tractable estimation by treating each aggregate as a primitive object of choice. The central question is whether choice frequencies over aggregates, generated by a RUM on the underlying alternatives, are also consistent with an ARUM. This paper addresses that question when the composition of each aggregate is heterogeneous across consumers, may vary across menus, and is unobservable to the analyst.
We characterize the testable implications of RU-rationality across three frameworks and show that they are substantially weaker than those of ARU-rationality. In the canonical outside-option environment, Theorem 3.1 establishes that RU-rationality is equivalent to Limited Monotonicity plus Partial RU-rationality. Theorem 3.2 gives a vertex characterization: the RU-rational set is a polytope with “menu-effect” deterministic vertices, and it strictly contains the ARU polytope. When the aggregation correspondence is fixed, Theorem 3.3 shows a strict nesting $\text{ARU}\subsetneq RU(2)\subsetneq\cdots\subsetneq RU(|\mathscr{A}_{A}|)$, with stabilization at $RU(|\mathscr{A}_{A}|+1)$, and Theorem 3.4 provides an approximation bound for $RU(n)$ below this threshold. All these results extend to the multiple aggregates case and imply that RU-rationality is far weaker than ARU-rationality.
We then identify two independent conditions under which the use of an ARUM is justified. Propositions 4.1–4.2 show that RU-rationality implies ARU-rationality under either non-overlapping preferences (within each aggregate, underlying alternatives are adjacent in every ranking in the support) or menu-independent composition. These results provide practical guidance on how to construct aggregated datasets in ways that justify the use of ARUM. Finally, simulations based on an underlying logit model quantify the empirical relevance of the gap: both estimation bias and the Euclidean distance to the ARU polytope increase as composition becomes more menu-dependent or preferences become more overlapping, and the resulting biases can be large enough to reverse inferred preference orderings and substantially distort relative odds ratios.
Appendix A Proofs
For ease of notation, we will assume throughout this section that each aggregate $a\in\mathscr{A}_{A}$ is composed of itself, that is $X(a)$ contains only $a$. Thus by abuse of notation, we may identify the atomic aggregates $\mathscr{A}_{A}$ and the corresponding set of underlying alternatives $\cup_{a\in\mathscr{A}_{A}}X(a)$.
A.1 Proof of Theorem 3.1 and Theorem 3.5
We provide a proof of Theorem 3.5, which implies Theorem 3.1 as a special case. [^25] We also prove Theorem 3.1 independently in Section D.1 to provide additional intuition for the first step of the construction. The alternate proof uses only $|\mathscr{A}_{N}|+1$ underlying alternatives belonging to $X(a_{0})$.
The necessity is trivial. We show the sufficiency. Fix an aggregation correspondence $X$ such that $|X(a)|=|\mathscr{A}_{A}|+2$ for all $a\in\mathscr{A}_{N}$. Such an aggregation correspondence exists by Richness of $\mathscr{X}$. For each $a\in\mathscr{A}_{N}$ and $y\in\mathscr{A}_{A}$, choose an element of $X(a)$, denoted by $x_{a}(y)$. We also fix two alternatives $\overline{x}_{a},\underline{x}_{a}\in X(a)$. Note that there exist such alternatives under the assumption that $|X(a)|=|\mathscr{A}_{A}|+2$ for all $a\in\mathscr{A}_{N}$. We again abuse notation and assume that each aggregate $a\in\mathscr{A}_{A}$ is composed of itself, that is $X(a)=\{a\}$.
By Partial RU-rationality, there exists a distribution $\tilde{\mu}$ over linear orders $\tilde{\succ}$ on $\mathscr{A}_{A}$ that rationalizes $\rho$ on subsets of $\mathscr{A}_{A}$. Based on this distribution, we will construct a RUM $\rho$ on the entire set $\mathscr{X}$ of the underlying alternatives as follows. For each ranking on $\mathscr{A}_{A}$ in the support of $\tilde{\mu}$, we construct ranking on $\mathscr{X}$: we place $x_{a}(y)$ right above $y$, $\overline{x}_{a}$ and $\underline{x}_{a}$ at the bottom with $\overline{x}_{a}$s above $\underline{x}_{a}$s. The order among non-atomic aggregates is arbitrarily fixed. For example, let $\{a_{0},a_{1}\}=\mathscr{A}_{N}$ and fix an order by index. Also let $\mathscr{A}_{A}=\{y_{0},y_{1}\}$ and $y_{0}\ \tilde{\succ}\ y_{1}$. Then, we define $\succ$ on $\mathscr{X}$ as follows:
Fix $D\subseteq\mathscr{A}_{A}$ and $E\subseteq\mathscr{A}_{N}$. In the following, we will construct $\lambda\in\prod_{b\in D\cup E}(2^{X(b)}\setminus\{\emptyset\})$ to establish the following equality for all $b\in D\cup E$:
The goal is to construct a probability distribution $\lambda$ on $\prod_{b\in E}(2^{X(b)}\setminus\{\emptyset\})$ that satisfies
and
The proof consists of two steps. At the first step, for each $a\in E$ we will construct a $\lambda^{a}$ that achieves (10) and puts the remaining choice probability on $a$. Given $\lambda^{a}$, in the second step, by taking a convex combination of $\lambda^{a}$, we will obtain the desired $\lambda$ that satisfy (10) and (11) for the given $D$ and $E$.
$=\rho(D\cup E,y_{0}).$
$=\rho(D\cup E,y_{n}),$
where the second equality holds because $x_{a}(y_{n})\succ y_{n}$ for all $\succ\in supp(\mu_{{\scriptscriptstyle\mathscr{X}}})$; the second to the last equality holds because $\sum_{(S_{b})_{b\in E}\in supp(\lambda^{a}_{n-1})\setminus(X(b))_{b\in E}}\lambda^{a}_{n-1}((S_{b})_{b\in E})=1-\lambda^{a}_{n-1}((X(b))_{b\in E})=1-\lambda^{a}_{0}((X(b))_{b\in E})=\lambda^{a}_{0}(\overline{x}_{a},(\underline{x})_{E\setminus a})$; and finally the third to the last equality holds because $y_{n}$ is a maximum element among $D$ with respect to $\succ$ if and only if $y_{n}$ is a maximum element among $D\cup(\cup_{b\in E}S_{b})$ with respect to $\succ$ (Note that this equivalence holds because $S_{b}\subseteq\{x_{b}(y_{i})|i<n\}\cup\{\underline{x}_{b},\overline{x}_{b}|b\in\mathscr{A}_{N}\}$ and $x_{b}(y_{i})$ is the immediate predecessor of $y_{i}$ for all $\succ\in supp(\mu_{{\scriptscriptstyle\mathscr{X}}})$). It follows that $\lambda^{a}$ satisfies (10) as desired: for all $y\in D$
Step 2: In the following, given $(\lambda^{a})_{a\in\mathscr{A}_{N}}$, we will construct $\lambda$ that satisfies (11) as well as (10). First observe that for all $a\in\mathscr{A}_{N}$,
Since $\sum_{a\in E}\rho(D\cup E,a)=1-\sum_{y\in D}\rho(D\cup E,y)$ and each $\lambda^{a}$ is a probability distribution on $\prod_{a\in E}2^{X(a)}\setminus\{\emptyset\}$, so is $\lambda$. Moreover, we have the desired equalities for choice frequency of each aggregate $b\in E$ as follows: